OpenStudy (anonymous):

A 10-ft diagonal brace on a bridge connects a support of the center of the bridge to a side support on the bridge. The horizontal distance that it spans is 2 ft longer that the height that it reaches on the side of the bridge. Find the horizontal and vertical distances spanned by this brace.

OpenStudy (anonymous):

so you know that for a right triangle, a^2+b^2=c^2, and c, is 10 here. you know that a=b+2, where a is the horizontal distance. Now you have two equations and two unknowns. Solve!